OpenStudy (ttp):

The face of a tent is shaped like a triangle with an area of 15.125ft^2 squared. The area of the triangle is represented by the formula A equals one half b times h. The height of the tent is 5.5 feet. How wide is the base of the tent? 5.5 feet 6.0 feet 6.5 feet 7.0 feet

OpenStudy (anonymous):

"The area of the triangle is represented by the formula A equals one half b times h." \[A = \frac{ 1 }{ 2 }b\times h\]plug in the values:\[(15.125) = \frac{ 1 }{ 2 }b\times (5.5)\]and then re-arrange the formula: \[(15.125) = \frac{ 1 }{ 2 }b\times (5.5)..... |:5.5\]\[\frac{(15.125)}{(5.5)} = \frac{ 1 }{ 2 }b............|\times 2\]